# AI-Enhanced 7-Point Checklist for Melanoma Detection Using Clinical Knowledge Graphs and Data-Driven Quantification

### v1

@@ 2.1 Clinical-Dermoscopic Multimodal Fusion (CD-MFM)
Step 1: Generate $\mathcal{G}_{\textrm{internal}}$ Internal Graph
Initialize an empty directed graph $\mathcal{G}_{\textrm{internal}}$;
Initialize a co-occurrence matrix $\mathcal{Q}$;
for each co-directional occurrence of nodes $p$ and $q$, denoted as $(p,q,c)$ , in $\mathcal{H}$ do
       Increment $\mathcal{Q}_{p\rightarrow q}$ by count $c$;
end for
for $p=1$ to $7$ do
       for $q=1$ to $7$ do
             if $\mathcal{Q}_{p\rightarrow q}\geq 1$ then
                   Calculate the edge weight as $w_{p\rightarrow q}=\frac{\mathcal{Q}_{p\rightarrow q}}{\text{total occurrencesof }v_{\textrm{PCL}_{p}}}$;
                   Add a directed edge from $v_{\textrm{PCL}_{p}}$ to $v_{\textrm{PCL}_{q}}$ with weight $w_{p\rightarrow q}$;
             end if
       end for
end for
Step 2: Generate $\mathcal{G}_{\textrm{external}}$ External Graph
Initialize an empty directed graph $\mathcal{G}_{\textrm{external}}$;
for $p=1$ to $7$ do
       if $\mathcal{Q}_{p,\textrm{Mel}}\geq 1$ then
             Calculate the edge weight $w_{\textrm{Mel}\rightarrow p}$ as $w_{\textrm{Mel}\rightarrow p}=\frac{\text{total occurrences of }p}{\text{totaloccurrences of Mel}}$;
             Add a directed edge from Mel to $v_{\textrm{PCL}_{p}}$ with weight $w_{\textrm{Mel}\rightarrow p}$;
       end if
      if $\mathcal{Q}_{\textrm{Mel},p}\geq 1$ then
             Calculate the edge weight $w_{p\rightarrow\textrm{Mel}}$ as $w_{p\rightarrow\textrm{Mel}}=\frac{\text{total occurrences of Mel}}{\text{total occurrences of }p}$;
             Add a directed edge from $v_{\textrm{PCL}_{p}}$ to Mel with weight $w_{p\rightarrow\textrm{Mel}}$;
       end if
end for
Step 3: Combine $\mathcal{G}_{\textrm{internal}}$ and $\mathcal{G}_{\textrm{external}}$
Calculate $\mathcal{G}=\alpha\cdot\mathcal{G}_{\textrm{internal}}+\beta\cdot\mathcal{G}_{\textrm{external}}$;
@@
2.2 7-Point Checklist Directed Graph Mining (7PCL-DirGM)
Within the 7PCL dataset, there exist both directional and causal relationships among the various attributes, each associated with different stages of melanoma development and their causes. For instance, both irregular pigment network and atypical pigmentation are important features to consider when evaluating potential melanomas. However, irregular pigment network is often considered a more specific and reliable indicator of melanoma when compared to atypical pigmentation alone. This is because the disruption of the normal pigment network structure is a hallmark feature of melanomas and is less commonly seen in benign moles. Therefore, the presence of an irregular pigment network raises the suspicion for melanoma and may prompt further evaluation or biopsy [de2018reinterpreting, pampena2020clinical, shi2020retrospective]. Recognizing these distinctions aids in filtering out superfluous and inaccurate interactions, enhancing the efficiency of graph network data utilization. Moreover, contrary to previous studies equating disease diagnosis labels with 7PCL features [kawahara2018seven, bi2020multi, tang2022fusionm4net, fu2022graph], we emphasize that the information pertaining to disease diagnosis, the various attributes, and interactions among these attributes represent significant directed information that should be considered separately with appropriate weighting for effective integration. To address these complexities, our study proposes the 7PCL-DirGM, a directional graph mining method tailored to extract vital attribute interaction data related to melanoma.
2.2.1 Graph Node Feature Encoding
As a critical input to graph convolutional learning networks, the encoding of node information is often overshadowed by the emphasis on topological relationships between nodes. Previous research has frequently relied on one-hot encoding to extract textual information features. However, this method tends to truncate the richness of information contained in node attributes, reducing them to mere labels. For example, both positive expressions related to pigmentation and streaks attributes tend to contain irregular descriptive statements, which is an important piece of information. In this study, we use Global Vectors (GloVe) [pennington2014glove], which can preserve semantic information, to extract node information features. Based on the predefined variables, $x^{i}_{\textrm{7p}_{j}}$ is transformed into ${X^{i}_{\textrm{7p}_{j}}\in\mathbb{R}^{7\times 128}}$, by utilizing the GloVe word embedding module, which employs word-content co-occurrence probabilities. Here, the length of encoded features is set to 128 based on pre-experiments.
2.2.2 Internal & External Directional Weighted Graph (IEDWG)
@@
2.2.3 Adaptive Receptive Field Proximity (ARFP)
When working with a weighted directed graph, the challenge initially lies in the absence of directed information along long paths, which results in insufficient global information and unbalanced receptive fields. Inspired by the digraph inception convolutional networks [tong2020digraph], we utilized the adjustable parameter $k$ to generate $k^{th}$-order proximity based on the information of both the meeting path and diffusion path. This approach enables the simultaneous extraction of direct and indirect connectivity information to the target and neighboring nodes, in contrast to relying solely on first- and second-order proximity. Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, for $k\geq 2$, there exists a node $e\in\mathcal{V}$, the meeting path and diffusion path are defined as:
$\mathcal{M}_{p,q}^{(k)}&=v_{p}\underbrace{\rightarrow\cdots\rightarrow}_{k-1\text{ edges }}v_{e}\underbrace{\leftarrow\cdots\leftarrow}_{k-1\text{ edges }}v_{q},\\ \mathcal{D}_{p,q}^{(k)}&=v_{p}\underbrace{\leftarrow\cdots\leftarrow}_{k-1\text{ edges }}v_{e}\underbrace{\rightarrow\cdots\rightarrow}_{k-1\text{ edges}}v_{q}.\end{aligned}\right.$ (4)
where $\mathcal{M}_{p,q}^{(k)}$ and $\mathcal{D}_{p,q}^{(k)}$ represent the meeting path and diffusion path between nodes $p$ and $q$$(p,q\in\mathcal{V})$. If both $\mathcal{M}_{p,q}{}^{(k)}$ and $\mathcal{D}_{p,q}{}^{(k)}$ exist between nodes $p$ and $q$, then we consider them to be at the $k^{\text{th}}$-order proximity level, with $v_{e}$ being their $k^{\text{th}}$-order common neighbor. It is worth noting that a node has $0^{\text{th}}$-order proximity with itself and $1^{\text{st}}$-order proximity with its directly connected neighbors.
Using the 7PCL directed weighted graph and the $k^{\text{th}}$-order proximity, the 7PCL $k^{\text{th}}$-order proximity matrix founded on the conventional spectral-based GCN method is denoted as:
@@ 2.2.3 Adaptive Receptive Field Proximity (ARFP)
where ${{A}}$ represents the adjacency matrix of the directed weighted graph ${\mathcal{G}}$, and ${{D}}$ represents the corresponding diagonalized degree matrix. The matrix intersection operation $\text{intersect}(\cdot)$ creates a new matrix with the element-wise intersection of both meeting and diffusion paths. This operation aims to symmetrize the $k^{th}$-order proximity matrix ${P}^{(k)}$ to fit the structure of spectral-based GCN. Here, the value of parameter $k$ represents the distance between two similar nodes, which determines the size of the receptive fields. By adjusting the value of parameter $k$, users can observe and select the optimal receptive field with scalable capabilities.
@@
where ${Z}^{(k)}$ represents the convolved output with dimensions $\mathbb{R}^{n\times d}$. ${X}$ denotes the node feature matrix with dimensions $\mathbb{R}^{n\times c}$. ${W}^{(k)}$ refers to the weight matrix, which is diagonalized from ${P}^{(k)}$, and ${\Theta}^{(k)}$, with dimensions $\mathbb{R}^{c\times d}$, is a trainable weight matrix. It’s noteworthy that when $k=1$, ${Z}^{(1)}$ is computed through digraph convolution using ${P}^{(1)}$ as defined in Eq. 5, and $\boldsymbol{\Pi}^{(1)}$ represents the corresponding approximate diagonalized eigenvector associated with ${P}^{(1)}$.
@@
Table 1: Classification Performance (AUC, Prec, Sens, and Spec) of the Proposed Method Compared with the State-of-the-Art Methods, Evaluated on the EDRA Dataset

### v3

@@ III-A Clinical-Dermoscopic Multimodal Fusion (CD-MFM)
2
3Step 1: Generate $\mathcal{G}_{\textrm{internal}}$ - Internal Graph
4Initialize an empty directed graph $\mathcal{G}_{\textrm{internal}}$;
5 Initialize a co-occurrence matrix $\mathcal{Q}$;
6for each co-directional occurrence $c$ of nodes $p$ and $q$, denoted as $(p,q,c)$ , in $\mathcal{H}$ do
7  Increment $\mathcal{Q}_{p\rightarrow q}$ by count $c$;
8
9 end for
10
11for $p=1$ to 77 do
12  for $q=1$ to 77 do
13     if $\mathcal{Q}_{p\rightarrow q}\geq 1$ then
14        Calculate the edge weight as $w_{p\rightarrow q}=\frac{\mathcal{Q}_{p\rightarrow q}}{\text{total occurrences of }v_{\textrm{PCL}_{p}}}$;
15        Add a directed edge from $v_{\textrm{PCL}_{p}}$ to $v_{\textrm{PCL}_{q}}$ with weight $w_{p\rightarrow q}$;
16
17     end if
18
19  end for
20
21 end for
22
23Step 2: Generate $\mathcal{G}_{\textrm{external}}$ - External Graph
24Initialize an empty directed graph $\mathcal{G}_{\textrm{external}}$;
25
26for $p=1$ to 77 do
27  if $\mathcal{Q}_{p,\textrm{Mel}}\geq 1$ then
28     Calculate the edge weight $w_{\textrm{Mel}\rightarrow p}$ as $w_{\textrm{Mel}\rightarrow p}=\frac{\text{total occurrences of }p}{\text{total occurrences of Mel}}$;
29     Add a directed edge from Mel to $v_{\textrm{PCL}_{p}}$ with weight $w_{\textrm{Mel}\rightarrow p}$;
30
31  end if
32 if $\mathcal{Q}_{\textrm{Mel},p}\geq 1$ then
33     Calculate the edge weight $w_{p\rightarrow\textrm{Mel}}$ as $w_{p\rightarrow\textrm{Mel}}=\frac{\text{total occurrences of Mel}}{\text{total occurrences of }p}$;
34     Add a directed edge from $v_{\textrm{PCL}_{p}}$ to Mel with weight $w_{p\rightarrow\textrm{Mel}}$;
35
36  end if
37
38 end for
39
40Step 3: Combine $\mathcal{G}_{\textrm{internal}}$ and $\mathcal{G}_{\textrm{external}}$
41Calculate $\mathcal{G}=\alpha\cdot\mathcal{G}_{\textrm{internal}}+\beta\cdot\mathcal{G}_{\textrm{external}}$;
42
@@ 34     Add a directed edge from $v_{\textrm{PCL}_{p}}$ to Mel with weight $w_{p\rightarrow\textrm{Mel}}$;
III-B 7-Point Checklist Directed Graph Mining (7PCL-DirGM)
Within the 7PCL dataset, there exist both directional and causal relationships among the various attributes, each associated with different stages of melanoma development and their causes. For instance, both irregular pigment network and atypical pigmentation are important features to consider when evaluating potential melanomas. However, irregular pigment network is often considered a more specific and reliable indicator of melanoma when compared to atypical pigmentation alone. This is because the disruption of the normal pigment network structure is a hallmark feature of melanomas and is less commonly seen in benign moles. Therefore, the presence of an irregular pigment network raises the suspicion for melanoma and may prompt further evaluation or biopsy [34, 35, 36]. Recognizing these distinctions aids in filtering out superfluous and inaccurate interactions, enhancing the efficiency of graph network data utilization. Moreover, contrary to previous studies equating disease diagnosis labels with 7PCL features [6, 19, 20, 22], we emphasize that the information pertaining to disease diagnosis, the various attributes, and interactions among these attributes represent significant directed information that should be considered separately with appropriate weighting for effective integration. To address these complexities, our study proposes the 7PCL-DirGM, a directional graph mining method tailored to extract vital attribute interaction data related to melanoma.
III-B1 Graph Node Feature Encoding
In graph convolutional learning networks, the encoding of node information—often critical—tends to be overshadowed by the emphasis on topological relationships between nodes. Previous research frequently relies on one-hot encoding to represent textual information features; however, this approach reduces the richness of node attributes to simple labels, limiting the depth of information available for learning. For instance, positive expressions related to attributes like pigmentation and streaks often include irregular descriptive terms, which carry significant semantic information.
In this study, we employ the BERT embedding strategy [37], which not only preserves semantic context but also dynamically captures node feature details. Each node feature, denoted as $x^{i}_{\textrm{7p}j}$, is transformed into an embedding $X^{i}_{\textrm{7p}_{j}}\in\mathbb{R}^{7\times 768}$} using the BERT module, where 768 to match the default output dimensionality of BERT. Here, 768 was chosen based on pre-experiments to optimize the representation of encoded features.
III-B2 Internal & External Directional Weighted Graph (IEDWG)
@@
(a) External connectivity
(b) Internal connectivity
(c) Combined connectivity
@@
III-B3 Adaptive Receptive Field Proximity (ARFP)
Accurately modeling attribute relationships is crucial for melanoma diagnosis using the 7PCL. However, conventional graph-based methods struggle to capture long-range dependencies, leading to imbalanced receptive fields and suboptimal feature learning. To address this, we introduce ARFP, which enhances the multi-order connectivity of the 7PCL-directed weighted graph, improving melanoma classification. Inspired by digraph inception convolutional networks [38], ARFP leverages the $k^{\text{th}}$-order proximity to incorporate both direct and indirect attribute interactions. Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents 7PCL attributes and $\mathcal{E}$ represents their directed relationships, the proximity paths are defined as:
@@ III-B3 Adaptive Receptive Field Proximity (ARFP)
where $\mathcal{M}_{p,q}^{(k)}$ and $\mathcal{D}_{p,q}^{(k)}$ represent the meeting and diffusion paths between attributes $p$ and $q$ $(p,q\in\mathcal{V})$. If both paths exist, the attributes are at the $k^{\text{th}}$-order proximity level, with $v_{e}$ as their common neighbor.
Using the 7PCL-directed weighted graph and the $k^{\text{th}}$-order proximity, the 7PCL $k^{\text{th}}$-order proximity matrix founded on the conventional spectral-based GCN method is denoted as:
[…]
where A{{A}} represents the adjacency matrix of the directed weighted graph $\mathcal{G}$}, and D{{D}} represents the corresponding diagonalized degree matrix. The matrix intersection operation $\text{intersect}(\cdot)$ creates a new matrix with the element-wise intersection of both meeting and diffusion paths. This operation aims to symmetrize the $k^{th}$-order proximity matrix ${P}^{(k)}$ to fit the structure of spectral-based GCN. Here, the value of parameter $k$ represents the distance between two attributes, which determines the size of the receptive fields. By adjusting the value of parameter $k$, users can observe and select the optimal receptive field with scalable capabilities.
@@
where ${Z}^{(k)}$ represents the convolved output with dimensions $\mathbb{R}^{n\times d}$. X{X} denotes the node feature matrix with dimensions $\mathbb{R}^{n\times c}$. ${W}^{(k)}$ refers to the weight matrix, which is diagonalized from ${P}^{(k)}$, and ${\Theta}^{(k)}$, with dimensions $\mathbb{R}^{c\times d}$, is a trainable weight matrix. It’s noteworthy that when $k=1$, ${Z}^{(1)}$ is computed through digraph convolution using ${P}^{(1)}$ as defined in Eq. 7, and $\bm{\Pi}^{(1)}$ represents the corresponding approximate diagonalized eigenvector associated with ${P}^{(1)}$.
This approach enables a scalable representation of 7PCL attribute relationships, ensuring that clinically meaningful dependencies are preserved and effectively utilized in melanoma diagnosis. By leveraging the adaptive receptive field, our framework refines multi-scale learning across directed attributes, improving classification performance and interpretability.
III-C Gradient Diagnostics with Data-Driven Weighting (GD-DDW)
In this module, we first establish a parallel multi-label classifier designed to address the classification of seven attributes on the checklist. As $X$ is the input, $y_{{\textrm{7p}}_{j}}=[y_{{\textrm{7p}}_{1}},y_{{\textrm{7p}}_{2}},...,y_{{\textrm{7p}}_{7}}]\in\{0,1\}^{7}$ is denoted as the ground truth of all seven attributes in the format of a binary indicator vector. In the proposed method, the focal loss function [39] was employed with the objective of enhancing the performance on the dataset that exhibiting a serious imbalance.
The focal loss function for each label can be defined as:
$L_{\textrm{7p}}=\sum_{j=1}^{7}L_{{\textrm{7p}}_{j}}$ (10)
$L={L_{\textrm{7p}}}+{\lambda L_{\textrm{mel}}}$ (12)
where $\lambda$ is the weighting parameter that adjusts the contributions of the two losses, specifically, making sure the loss values are in the same order of magnitude.
IV Experiments
IV-A Materials
In this study, we utilized the publicly available EDRA dataset (also known as Derm7pt), specifically curated for 7PCL studies and annotated by Kawahara et al. [6]. This dataset comprises paired dermoscopic and clinical images sourced from 1011 patients, each image has a maximum resolution of 768 by 512 pixels. Notably, nine patients lacked clinical images, which were substituted with corresponding dermoscopic ones. Throughout our study, we strictly adhered to the dataset’s official specifications, which stipulated a total of 413 training, 203 validation, and 395 test images. Concerning the labels, the 7PCL attributes were incorporated into the analysis, following the official categorization. It is worth noting that the original disease classification (DIAG) system included specific disease labels for MISC, BCC, Mel, SK, and Nevi. This study primarily focused on distinguishing melanoma from lookalike diseases. To verify the generalizability of our proposed method, we also used two publicly available datasets, ISIC 2017 and ISIC 2018 [17], which also have multiple attributes. The visual attributes contained in ISIC 2017 include milia-like cysts, negative network, pigment network, and streaks. A total of 2000 training, 150 validation, and 600 test images. ISIC 2018 adds an additional attribute, globules, to the 2017 version and adds a portion of the data to make the data distribution adjusted to 2594 training, 100 validation, and 1000 test images.
IV-B Implementation Details
In our study, we ensured the integrity of our results by employing consistent data processing methods and adhering to uniform model training protocols across various ablation comparison experiments. Our chosen deep learning model framework is based on three residual blocks. The entire architecture is meticulously fine-tuned end-to-end, utilizing the Adam optimizer [40] with a concurrent learning rate set to 0.00001. We predefined the number of epochs for pretraining to 150, with an early-stopping mechanism in place. If the model’s performance on the validation set fails to improve over 50 consecutive epochs, the training process halts, and the best-performing model is saved. The loss functions utilized are based on the focal loss to circumvent the result bias that may arise from imbalanced data. The hyperparameters for the different stages outlined in the methodology were selected based on insights gained from the training process and represent approximate ranges. The initial stage of the process is the pre-processing of the images. To ensure consistency and minimize the impact of varying image formats, we removed black edges from the images and resized them to 512 by 512 pixels, maximizing the use of image features while avoiding information loss.
During the data preparation phase, all labels are transformed into standardized textual descriptions. These are then processed by the semantic embedding function to effectively utilize label information. Each patient case follows the format: ” This patient has been diagnosed with [Diagnosis], exhibiting the attributes [Attribute 1], [Attribute 2], and [Attribute 3]. ” This structured representation ensures consistency in data processing and facilitates the meaningful integration of clinical attributes into the model. The standard texts of the indicated 7PCL are atypical pigment network (ATP-PN), irregular streaks (IR-STR), irregular pigmentation (IR-PIG), regression structure (RS), irregular dots and globus (IR-DaG), blue-whitish veil (BWV), and irregular vascular structures (IR-VS).
In the CD-MFM section, features from different channels undergo two stages of fusion using weighted averaging. Initially, a fusion process occurs between clinical and dermoscopy images using the dual-attention blocks, followed by a second fusion stage after incorporating topological relations learned by the graph network. The parameters $\gamma$ series in Eq. 5 are configured to 1/3{1}/{3} to mitigate potential biases stemming from assigning varying weights to the same feature across multiple instances. In constructing the graph network, we consider only a subset of the possible connections between nodes due to the directionality of the multistep connections. Based on the average number of co-existing labels in the training set and the avoidance of isolated points, the connections are ordered according to the edge weights, ensuring that each node has at least one and no more than three edges. The order of connection $k$ is set to 3 due to the natural limitation of the number of nodes in 7PCL. In the final diagnostic phase, the parameter $\lambda$, utilized in the global optimization of the loss function in Eq. 12, is set to 1, given that both loss values are of a comparable magnitude. All experiments were conducted on one NVIDIA Tesla V100 GPU card (32 GB memory). The duration of each training cycle is approximately one hour.
IV-C Evaluation Settings
The study used standard performance metrics such as AUC, sensitivity and specificity. To demonstrate our method’s advantages, we compared it with state-of-the-art methods like RemixFormer++ [31], CAFNet [30], GRM-DC [22], 7PCL-Constrained-CC [25], HcCNN [19], TripleNet [41], CCRD [42], GELN [43], and EmbeddingNet [44]. Specifically, using the Inception models introduced by Kawahara et al. [6] as the baseline for comparison, as they originally published the data along with the method.
We also conducted ablation experiments to evaluate the critical components of our method. These experiments included comparisons between unimodal and multimodal fusion (Dual-Attention), assessments of our directional graph mining-based GCN strategy (7PCL-DirGM), and tests of our gradient-driven prediction structure (GD-DDW) against parallel architectures.
In addition, the prediction results for all attributes obtained from the first layer of the GD-DDW structure were utilized. Following the traditional 7PCL rule, diagnostic scores were calculated by summing the assigned attribute scores. Thresholds ranging from 1 to 7 were applied to diagnose melanoma. These results were then compared with the proposed method across two tasks: Mel vs. Nevi and Mel vs. other diseases, including MISC, SK, Nevi, and BCC.
We further validated our method using two additional publicly available datasets, ISIC 2017 and ISIC 2018. Although these datasets provide only attribute information and lack diagnostic labels, preventing the evaluation of GD-DDW’s impact, they still allow the validation of directed multi-order relationships between the attributes (CKTG).
@@ IV-C Evaluation Settings
TABLE I: Classification Performance (AUC, Sens, and Spec) of the Proposed Method Compared with the State-of-the-Art Methods, Evaluated on the EDRA Dataset. Top two AUC scores per label are highlighted.